# OCMDP: Observation-Constrained Markov Decision Process

### v1

@@ 5.2 Experiments on the HeartPole Healthcare Simulator
In this study, we employ the HearPole environment [17] for our experimental evaluations [^1] https://github.com/vadim0x60/heartpole. HearPole is a simplified, rule-based healthcare simulation developed to facilitate the benchmarking of reinforcement learning (RL) algorithms in medical contexts. Designed as an OpenAI Gym environment, HearPole serves as an introductory platform for testing RL techniques before applying them to more complex clinical datasets.
HearPole models a healthcare scenario where the agent’s objective is to maximize productivity while maintaining the patient’s health. The action space consists of four discrete actions: do nothing, drink coffee, drink beer, and sleep. These actions influence the patient’s state, which is represented by six continuous observation variables: alertness, hypertension, intoxication, time since last sleep, time elapsed, and work done. The agent receives a reward of +1 for each unit of work completed, incentivizing productivity, and incurs a penalty of -100 if a heart attack occurs, emphasizing the importance of maintaining patient health. Each dimensional observation incurs a cost of $C_{O_{i}}=-0.05$. Each episode is capped at 1000 steps to balance sufficient exploration and training efficiency.
The controlled and transparent nature of HearPole allows researchers to systematically evaluate and refine RL algorithms within a healthcare-themed setting. By providing a clear and manageable environment, HearPole enables the assessment of various RL methodologies, including policy-based and value-based approaches, in achieving a balance between productivity and health maintenance. This setup is particularly advantageous for identifying strengths and limitations of different algorithms in a simplified yet relevant medical scenario before transitioning to more intricate and data-intensive environments.
Through the use of HearPole, we can effectively benchmark our proposed RL algorithms against established baseline methods. The discrete action space and well-defined reward structure facilitate a focused analysis of algorithm performance, stability, and convergence. Additionally, the simulation’s ability to generate reproducible and interpretable results supports a thorough investigation into the decision-making processes of the agents, providing insights that are transferable to real-world clinical applications.
@@ 5.3 Experimental Results
Figures 5 and 6 illustrate the performance of our proposed Observation-Constrained MDP (OCMDP) framework across two tasks: the Diagnostic Chain Task and the HearPole Task. The results demonstrate significant improvements in both control performance and observation efficiency, validating the versatility and effectiveness of OCMDP across different complex environments. In the Diagnostic Chain Task (Figure 5), our method achieves a relative improvement of 71% in expected cumulative reward over the baseline model-free control policies. This improvement reflects the effectiveness of our control policy optimization in maximizing rewards while accounting for observation costs. Furthermore, the optimization of the observation policy reduces observation costs by 50% compared to fixed and continuous observation strategies, as shown in Figure 5.(b), indicating that OCMDP successfully balances the trade-off between acquiring valuable information and minimizing associated costs.
In the Heartpole Task (Figure 6), OCMDP outperforms several baseline reinforcement learning algorithms, with a relative improvement of approximately 75% in episodic return over the next best-performing algorithm, PPO. This improvement highlights OCMDP’s ability to effectively balance observation costs with control rewards, particularly in diagnostic and decision-making tasks with high complexity. Figure 6.(b) compares different observation strategies within OCMDP, including optimal observation (our adaptive strategy), always observe, and never observe. The optimal observation strategy yields the highest episodic return, achieving approximately 80% higher return compared to the always observe setting, while incurring fewer observation costs. Additionally, it outperforms the never observe setting by a relative improvement of 90% in episodic return, underscoring the importance of selective observation to enhance control performance on the HearPole Task.

### v4

@@ 5.2 Experiments on the HeartPole Healthcare Simulator
In this study, we employ the HeartPole environment [17] for our experimental evaluations [^1] https://github.com/vadim0x60/heartpole. HeartPole is a simplified, rule-based healthcare simulation developed to facilitate the benchmarking of reinforcement learning (RL) algorithms in medical contexts. Designed as an OpenAI Gym environment, HeartPole serves as an introductory platform for testing RL techniques before applying them to more complex clinical datasets.
HeartPole models a healthcare scenario where the agent’s objective is to maximize productivity while maintaining the patient’s health. The action space consists of four discrete actions: do nothing, drink coffee, drink beer, and sleep. These actions influence the patient’s state, which is represented by six continuous observation variables: alertness, hypertension, intoxication, time since last sleep, time elapsed, and work done. The agent receives a reward of +1 for each unit of work completed, incentivizing productivity, and incurs a penalty of -100 if a heart attack occurs, emphasizing the importance of maintaining patient health. Each dimensional observation incurs a cost of $C_{O_{i}}=-0.05$. Each episode is capped at 1000 steps to balance sufficient exploration and training efficiency.
The controlled and transparent nature of HeartPole allows researchers to systematically evaluate and refine RL algorithms within a healthcare-themed setting. By providing a clear and manageable environment, HeartPole enables the assessment of various RL methodologies, including policy-based and value-based approaches, in achieving a balance between productivity and health maintenance. This setup is particularly advantageous for identifying strengths and limitations of different algorithms in a simplified yet relevant medical scenario before transitioning to more intricate and data-intensive environments.
Through the use of HeartPole, we can effectively benchmark our proposed RL algorithms against established baseline methods. The discrete action space and well-defined reward structure facilitate a focused analysis of algorithm performance, stability, and convergence. Additionally, the simulation’s ability to generate reproducible and interpretable results supports a thorough investigation into the decision-making processes of the agents, providing insights that are transferable to real-world clinical applications.
@@ 5.3 Experimental Results
Figures 5 and 6 illustrate the performance of our proposed Observation-Constrained MDP (OCMDP) framework across two tasks: the Diagnostic Chain Task and the HeartPole Task. The results demonstrate significant improvements in both control performance and observation efficiency, validating the versatility and effectiveness of OCMDP across different complex environments. In the Diagnostic Chain Task (Figure 5), our method achieves a relative improvement of 71% in expected cumulative reward over the baseline model-free control policies. This improvement reflects the effectiveness of our control policy optimization in maximizing rewards while accounting for observation costs. Furthermore, the optimization of the observation policy reduces observation costs by 50% compared to fixed and continuous observation strategies, as shown in Figure 5.(b), indicating that OCMDP successfully balances the trade-off between acquiring valuable information and minimizing associated costs.
In the Heartpole Task (Figure 6), OCMDP outperforms several baseline reinforcement learning algorithms, with a relative improvement of approximately 75% in episodic return over the next best-performing algorithm, PPO. This improvement highlights OCMDP’s ability to effectively balance observation costs with control rewards, particularly in diagnostic and decision-making tasks with high complexity. Figure 6.(b) compares different observation strategies within OCMDP, including optimal observation (our adaptive strategy), always observe, and never observe. The optimal observation strategy yields the highest episodic return, achieving approximately 80% higher return compared to the always observe setting, while incurring fewer observation costs. Additionally, it outperforms the never observe setting by a relative improvement of 90% in episodic return, underscoring the importance of selective observation to enhance control performance on the HeartPole Task.